2. A firm sells its product in a perfectly competitive market where other firms charge a price of $80 per unit. The firms total costs are C(O) 40 80202 a. How much output should the firm produce in the short run? b. What price should the firm charge in the short run? c. What are the firms short-run profits? d. What adjustments should be anticipated in the long run?
Answer #1

a) MC = 8 + 4Q

The profit maximization condition is:

MC = P

8 + 4Q = 80

4Q = 80 - 8 = 72

Q = 72 / 4 = 18

Thus, the firm should produce 18 units in the short run.

b) The firm should charge $80 in the short run. Because in a competitive market a firm is a price taker. Since other firms charge $80, this firm will also charge $80.

c) TR = P * Q = $80 * 18 = $1,440

    TC = 40 + 8Q + 2Q2 = 40 + 8(18) + 2(18)2 = 40 + 144 + 648 = $832

     Profit = TR - TC = $1,440 - $832 = $608

d) Since the firms are earning positive economic profit in the short run, therefore, new firms will enter into the industry in the long run. This will lead to an increase in supply of the good which in turn leads to decrease in equilibrium price level. This entry of new firms will continue till all the firms earns normal profit.